OpenStudy (anonymous):

19 is 10 percent of what number

myininaya (myininaya):

\[19=10 \% n\]

myininaya (myininaya):

Solve for n

OpenStudy (fj):

190

myininaya (myininaya):

\[19=\frac{10}{100}n\] \[19=\frac{1}{10}n\] \[19 \cdot 10=n\] \[n=190\]